# MP5: A Multi-modal Open-ended Embodied System in Minecraft via Active Perception

### v3

@@ Abstract
It is a long-lasting goal to design an embodied system that can solve long-horizon open-world tasks in human-like ways. However, existing approaches usually struggle with compound difficulties caused by the logic-aware decomposition and context-aware execution of these tasks. To this end, we introduce MP5, an open-ended multimodal embodied system built upon the challenging Minecraft simulator, which can decompose feasible sub-objectives, design sophisticated situation-aware plans, and perform embodied action control, with frequent communication with a goal-conditioned active perception scheme. Specifically, MP5 is developed on top of recent advances in Multimodal Large Language Models (MLLMs), and the system is modulated into functional modules that can be scheduled and collaborated to ultimately solve pre-defined context- and process-dependent tasks. Extensive experiments prove that MP5 can achieve a $22\%$ success rate on difficult process-dependent tasks and a $91\%$ success rate on tasks that heavily depend on the context. Moreover, MP5 exhibits a remarkable ability to address many open-ended tasks that are entirely novel.
Figure 1: The process of finishing the task “kill a pig with a stone sward during the daytime near the water with grass next to it.”. (a) To achieve the final goal (i.e., $\mathcal{O}_{8}$: “kill a pig Refer to caption”), a player should accomplish a list of sub-objectives $\{\mathcal{O}_{i}\}_{i=1}^{7}$ sequentially. During this process, the player should also be aware of some items in the environment, e.g., “grass Refer to caption”, “day Refer to caption” and etc. (b) This diagram shows the number of these necessary items in the context that should be perceived for each sub-objective, during the task execution process. (c) Images marked by $\mathcal{O}_{1}$ and $\mathcal{O}_{6}$ show the observed ego-centric views in the process of achieving the corresponding sub-objectives. Images marked by $\mathcal{O}_{8}^{1}$ and $\mathcal{O}_{8}^{2}$ indicate how the player executes the action about the last sub-objective “kill a pig Refer to caption”. This exemplar process tells that such long-horizon open-world embodied tasks in Minecraft should be solved both in the process-dependent and context-dependent way.
footnotetext: $*{}^{*}$ Equal contribution    ${}^{\dagger}$ Corresponding author   ${}^{\ddagger}$ Project leader
@@ 2.1 Multi-modal Large Language Models
With the development of Large Language Models (LLMs) like the GPT series [28, 2, 26], as well as open-source LLMs such as the LLaMA series [31, 32] and Vicuna [5], Multi-modal Large Language Models (MLLMs) have emerged. Examples of such MLLMs include LLaVA [20], InstructBLIP [6], and LAMM [38], among others [40, 37, 10, 4, 27]. In this work, we introduce an MLLM called MineLLM, which is specifically designed and trained for Minecraft, and leverage its perception, interaction, and analysis capabilities to build up Percipient for MP5, and further enable an objective-conditioned active perception scheme.
@@ 2.2 Agents in Minecraft
Previous works[11, 9, 39, 3, 18, 7, 21] attempt to use approaches such as hierarchical RL, goal-based RL, and reward shaping to train an agent in Minecraft. MineCLIP [9] enables the resolution of various open-ended tasks specified in free language, even without any manually designed dense rewards. DreamerV3 [12] succeeds in training agents in Minecraft with a learned world model. VPT [1] builds a foundation model for Minecraft by learning from massive videos. Based on VPT, Steve-1 [17] also explores bringing in MineCLIP [9] to get an instruction following policy with high performance. The development of recent large language model-related work Voyager [33], DEPS [35], GITM [41] further promote the advancement of agents in long-horizon tasks. These works use pre-trained large language models as the zero-shot planners[14] for agents, leveraging the powerful reasoning capabilities of large language models to obtain continuous operation instructions or executable policy lists.
@@ 3.1 Overview
As demonstrated in Fig. 2, our MP5 includes five major modules, i.e., Parser, Percipient, Planner, Performer, and Patroller. To be specific, Percipient is a parameter-efficiently fine-tuned Multimodal Large Language Model (MLLM) that is specified to the Minecraft environment. The Parser, Planner, and Patroller are pre-trained Large-language Models (LLMs). We also include retrieval-augmented generation (RAG) to enhance the quality of responses generated by Parser and Planner. Performer is an interface that explains each action from the action sequence into executable commands that directly control the game character.
@@
Why can MP5 solve context-dependent and process-dependent tasks? MP5 includes an active perception scheme by means of multi-round interactions between Percipient and Patroller, which is to actively perceive the environmental information in the observed images, with respect to various objectives raised by Planner or Performer. With the help of this scheme, Planner can schedule or update action sequences in awareness of the observed images, inventory status and etc., resulting in a situation-aware planning; Performer can execute actions that are adapted to the embodied environment, resulting in a embodied action execution. Patroller in this scheme can also feedback on better choices of plans/actions based on the visual evidence so that the process-dependent tasks are solved with fewer chances of context-dependent execution failures. Moreover, Percipient can understand open-ended visual concepts, therefore it allows MP5 to solve tasks that are never seen before. Based on these merits, MP5 is able to solve context-dependent and process-dependent tasks in an open-world embodied environment.
How does MP5 function? In Fig. 2, upon receiving a high-level task, MP5 first utilizes the Parser to generate a sequence of short-horizon sub-objectives, as a list of rich instructions in natural languages. The feasibility of the generated sub-objectives is augmented by retrieving an external Knowledge Memory (e.g., online wiki, user tips, and etc.). To one sub-objective, Planner schedules the action sequence that is grounded by the environmental information gathered by the active perception scheme. In this case, Performer will execute the actual actions by explaining the action sequence that is adapted to the embodied environment, via frequent interaction with the active perception scheme. Once there are execution failures (determined by Patroller), Planner will re-schedule the action sequence of the current sub-objective, or even update the following sub-objectives if some necessary sub-objectives are missing. Otherwise, the agent will go to the next sub-objective and schedule new action sequences, whilst the successful action sequence of the current sub-objective will be stored in the external memory of Planner (called Performer Memory), along with the agent situation when it was planned. In the end, the agent will stop when the final sub-objective of the task has been reached.
@@ 3.2 Active Perception
Similarly, active perception used in situation-aware planning is similar to what is explained here, except that the applied instructions do not contain the executable action. For more details please check the Sup. D.
@@ 3.3 Perception-aware Planning and Execution
Embodied Action Perception. As indicated in Sec. 3.2, Performer would like to communicate with the active perception scheme in every round of action execution, so as to enhance the ego-centric awareness of the agent. The new action will be executed if Patroller identifies necessary environmental information in the observed images that matches both the sub-objective and the goal of the current action. Otherwise, the current action is kept executing until encountering execution failures or the end of the episode. The successful action sequence about one sub-objective will be stored in the Performer Memory, together with necessary situational information of the agent when it was planned. For more details about the planning and execution process, please check Sup. F.2 and Sup. A.2.
@@ 3.4 Implementation Details
Percipient. The network of Percipient is depicted in Fig. 4. Images are processed by a frozen vision encoder MineCLIP [9], whose features are projected by an Alignment Net(we use two-layer MLP like LLaVA-1.5 [19]) to the same feature space as the text embeddings of the applied LLM (we use Vicuna-13B-v1.5 [5]). Then the vision and text tokens are concatenated to feed into a LoRA-based fine-tuned LLM [13]. We add LoRA [13] parameters to all projection layers of the self-attention layers in the LLM. Only the parameters of the Alignment Net and the LoRA module are optimized during training. The construction of the training data with respect to Percipient is in the Sup. A.1.
Parser, Planner, and Patroller. We utilize OpenAI’s GPT-4 [25] as LLMs in Parser, Patroller, and Planner. We also evaluate other alternatives of GPT-4 [25], such as open-source models like Vicuna-13B-v1.5 [5] and LLaMA2-70B-Chat [32] in Sup.C.3.
Performer. It is important to clarify that the actions generated by Planner are not low-level commands such as keyboard and mouse operations [1], but a set of simple actions (such as equip, move, craft). Inspired by GITM [41], we implement these actions appropriately through basic operations provided by the MineDojo [9] simulator. For more details, please check the Sup. A.2.
@@ 4.1 Experimental Setup
4.1.1 Environment Setting
We employ MineDojo [9] as the Minecraft simulation environment to build and validate MP5. We capture player ego-view images provided by MineDojo [9] as input of MP5, and further construct a dataset for training MineLLM. As for the output of MP5, we encapsulate MineDojo’s [9] actions to create our own action space.
4.1.2 Task Setting
To evaluate how our MP5 can organically integrate perception information with planning and execution, we mainly define two types of tasks: Context-Dependent Tasks and Process-Dependent Tasks as illustrated in Tab. 1 and Tab. 2.
Context-Dependent Tasks. primarily study how Active Perception enables the agent to better perceive low-level context information in the environment. We first establish $6$ aspects of environmental information derived from the Minecraft game environment: [Object, Mob, Ecology, Time, Weather, Brightness]. Each aspect has multiple options. For example, pigs [Uncaptioned image], cows [Uncaptioned image], and sheep [Uncaptioned image] are all elements belonging to Mob. Based on this, we define $16$ tasks and organize their difficulty into four levels by taking into account the number of information elements that require perception, as is shown in Tab. 1. For example, Easy tasks necessitate the perception of only one element, whereas Complex tasks involve the perception of $4$ to $6$ elements. We rigorously assess MP5’s proficiency in environmental context perception across these $16$ tasks. In Context-Dependent Tasks, our environment details are predetermined (e.g., biomes [Uncaptioned image], weather [Uncaptioned image], and etc.), as certain targets are exclusive to specific environments. Without this environmental specificity, the agent might never encounter the intended target. We retain each observation of active perception throughout the task, using them as references to ascertain the agent’s successful completion of the task.
Process-Dependent Tasks. focus on exploring the contributions of situation-aware planning, embodied action execution, and the integration with Active Perception in accomplishing long-term tasks while constantly perceiving the environment and dynamically adjusting actions. We select $25$ tasks from the technology tree and define their difficulty levels as Basic level [Uncaptioned image] to Diamond level [Uncaptioned image] based on the number of reasoning steps required to complete the tasks. All environmental factors (e.g., biomes [Uncaptioned image], weather [Uncaptioned image], and etc.) are randomized in Process-Dependent Tasks. More details can be found in Sup.C.1.
@@ 4.1.2 Task Setting
4.1.3 Evaluation Metrics
For different tasks, the agent’s initial position and environment seed are randomized. The agent begins in survival mode, commencing with an empty inventory, and faces the challenge of hostile mob generation. It starts from scratch, with a game time limit of 10 minutes, a time period equivalent to 12,000 steps at a control frequency of 20Hz. More details can be found in Sup. B.
@@ 4.1.3 Evaluation Metrics
In practice, we conduct $50$ games on Context-Dependent Tasks and $30$ games on Process-Dependent Tasks, averaging the success rates for both. The results are grouped according to the previously defined difficulty levels, and report the group means. For detailed definitions of the evaluation, please refer to Sup. C.
@@ 4.2.1 Results of Context-Dependent Tasks
In Context-Dependent Tasks, we primarily investigate how to enhance an agent’s perception of context information within the environment. We demonstrate the performance difference between Active Perception and other perception methods. We compare them with pre-trained multi-modal large language models LLaVA-1.5 [19] and GPT-4V [24], and analyze the performance of both active and fine-grained global perception on the tasks in Tab. 3. Although fine-grained global perception can obtain comprehensive perceptual information, due to the lack of objective-conditioned attention, the objective-related information obtained may be lacking or incorrect. Active perception only focuses on objective-related information and ignores other useless information, so that more accurate objective-related information can be obtained and better performance in Context-Dependent Tasks can be achieved. For the comparison, we use MineLLM, which is fine-tuned on the Minecraft instruction dataset we collect, slightly better than GPT-4V [24], which is trained on massive data, and substantially better than LLaVA-1.5 [19], which is not fine-tuned on instruction data. The complete results of Context-Dependent Tasks can be found in Sup.C.2.
@@ 4.2.2 Results of Process-Dependent Tasks
In Process-Dependent Tasks, we report the performance of the agent in completing long-horizon tasks by continuously perceiving the environment context and dynamically adjusting its actions. We also investigate the agent’s behavior in scenarios of non-situation-aware planning and non-embodied action execution. The complete results of Process-Dependent Tasks can be found in Sup.C.2.
@@ 4.2.3 Open-Ended Tasks
Processing long-horizon reasoning and understanding complex contexts are interconnected in the real world. For simplicity and comparability of the experimental setup, the first two task settings do not consider the intersection of process and context, as we cannot exhaust all combinations that these two task dimensions can form. Therefore, we refer to tasks that incorporate both Process-Dependent and Context-Dependent elements as Open-Ended Tasks. Specifically, these tasks require the agent to perceive different information of the environment at multiple stages of completing sub-objectives. As shown in Fig. 5, we present an example of an Open-Ended Task, named “Dig a block of sand [Uncaptioned image] near the water [Uncaptioned image] at night [Uncaptioned image] with a wooden shovel [Uncaptioned image]. We conduct extensive validations on this type of task, proving that MP5 can complete long-sequential tasks in challenging environments. More demonstrations and experimental results of Open-Ended Tasks can be found in Sup.E.3.

### v4

@@ Abstract
It is a long-lasting goal to design an embodied system that can solve long-horizon open-world tasks in human-like ways. However, existing approaches usually struggle with compound difficulties caused by the logic-aware decomposition and context-aware execution of these tasks. To this end, we introduce MP5, an open-ended multimodal embodied system built upon the challenging Minecraft simulator, which can decompose feasible sub-objectives, design sophisticated situation-aware plans, and perform embodied action control, with frequent communication with a goal-conditioned active perception scheme. Specifically, MP5 is developed on top of recent advances in Multimodal Large Language Models (MLLMs), and the system is modulated into functional modules that can be scheduled and collaborated to ultimately solve pre-defined context- and process-dependent tasks. Extensive experiments prove that MP5 can achieve a $22\%$ success rate on difficult process-dependent tasks and a $91\%$ success rate on tasks that heavily depend on the context. Moreover, MP5 exhibits a remarkable ability to address many open-ended tasks that are entirely novel. Please see the project page at https://iranqin.github.io/MP5.github.io/.
Figure 1: The process of finishing the task “kill a pig with a stone sward during the daytime near the water with grass next to it.”. (a) To achieve the final goal (i.e., $\mathcal{O}_{8}$: “kill a pig Refer to caption”), a player should accomplish a list of sub-objectives $\{\mathcal{O}_{i}\}_{i=1}^{7}$ sequentially. During this process, the player should also be aware of some items in the environment, e.g., “grass Refer to caption”, “day Refer to caption” and etc. (b) This diagram shows the number of these necessary items in the context that should be perceived for each sub-objective, during the task execution process. (c) Images marked by $\mathcal{O}_{1}$ and $\mathcal{O}_{6}$ show the observed ego-centric views in the process of achieving the corresponding sub-objectives. Images marked by $\mathcal{O}_{8}^{1}$ and $\mathcal{O}_{8}^{2}$ indicate how the player executes the action about the last sub-objective “kill a pig Refer to caption”. This exemplar process tells that such long-horizon open-world embodied tasks in Minecraft should be solved both in process-dependent and context-dependent way.
footnotetext: Equal contribution     Corresponding author   Project leader
@@ 2.1 Multi-modal Large Language Models
With the development of Large Language Models (LLMs) like the GPT series [29, 2, 27], as well as open-source LLMs such as the LLaMA series [32, 33] and Vicuna [5], Multi-modal Large Language Models (MLLMs) have emerged. Examples of such MLLMs include LLaVA [21], InstructBLIP [6], and LAMM [39], among others [42, 38, 10, 4, 28, 15]. In this work, we introduce MineLLM, which is specifically designed and trained for Minecraft, and leverage its perception, interaction, and analysis capabilities to build up Percipient for MP5, and further enable an objective-conditioned active perception scheme.
@@ 2.2 Agents in Minecraft
Previous works[11, 9, 40, 3, 19, 7, 22, 41] attempt to use approaches such as hierarchical RL, goal-based RL, and reward shaping to train an agent in Minecraft. MineCLIP [9] enables the resolution of various open-ended tasks specified in free language, even without any manually designed dense rewards. DreamerV3 [12] succeeds in training agents in Minecraft with a learned world model. VPT [1] builds a foundation model for Minecraft by learning from massive videos. Based on VPT, Steve-1 [18] also explores bringing in MineCLIP [9] to get an instruction following policy with high performance. The development of recent large language model-related work Voyager [34], DEPS [36], GITM [43] further promote the advancement of agents in long-horizon tasks. These works use pre-trained large language models as the zero-shot planners[14] for agents, leveraging the powerful reasoning capabilities of large language models to obtain continuous operation instructions or executable policy lists.
@@ 3.1 Overview
As demonstrated in Fig. 2, our MP5 includes five major modules, i.e., Parser, Percipient, Planner, Performer, and Patroller. Specifically, Percipient is a parameter-efficiently fine-tuned Multimodal Large Language Model (MLLM) that is specified to the Minecraft environment. The Parser, Planner, and Patroller are pre-trained Large-language Models (LLMs). We also include retrieval-augmented generation (RAG) to enhance the quality of responses generated by Parser and Planner. Performer is an interface that explains each action from the action sequence into executable commands that directly control the game character.
@@
Why can MP5 solve context-dependent and process-dependent tasks? MP5 includes an active perception scheme by means of multi-round interactions between Percipient and Patroller, which is to actively perceive the environmental information in the observed images, with respect to various objectives raised by Planner or Performer. With the help of this scheme, Planner can schedule or update action sequences in awareness of the observed images, inventory status and etc., resulting in a situation-aware planning; Performer can execute actions that are adapted to the embodied environment, resulting in a embodied action execution. Patroller in this scheme can also feedback on better choices of plans/actions based on the visual evidence so that the process-dependent tasks are solved with fewer chances of context-dependent execution failures. Moreover, Percipient can understand open-ended visual concepts, therefore it allows MP5 to solve tasks that are never seen before.
How does MP5 function? In Fig. 2, upon receiving a high-level task, MP5 first utilizes the Parser to generate a sequence of short-horizon sub-objectives, as a list of rich instructions in natural languages. The feasibility of the generated sub-objectives is augmented by retrieving an external Knowledge Memory. This knowledge mainly comes from three sources: part of it is from the online wiki, another part is from the crafting recipes of items in MineDojo [9], and some are user tips from Reddit. To one sub-objective, Planner schedules the action sequence that is grounded by the environmental information gathered by the active perception scheme. In this case, Performer will execute the actual actions by explaining the action sequence that is adapted to the embodied environment, via frequent interaction with the active perception scheme. Once there are execution failures (determined by Patroller), Planner will re-schedule the action sequence of the current sub-objective, or even update the following sub-objectives if some necessary sub-objectives are missing. Otherwise, the agent will go to the next sub-objective and schedule new action sequences, whilst the successful action sequence of the current sub-objective will be stored in the external memory of Planner (called Performer Memory), along with the agent situation when it was planned. In the end, the agent will stop when the final sub-objective of the task has been reached.
@@ 3.2 Active Perception
Similarly, active perception used in situation-aware planning is similar to what is explained here, except that the applied instructions do not contain the executable action. For more details please check the Sup. E.
@@ 3.3 Perception-aware Planning and Execution
Embodied Action Perception. As indicated in Sec. 3.2, Performer would like to communicate with the active perception scheme in every round of action execution, so as to enhance the ego-centric awareness of the agent. The new action will be executed if Patroller identifies necessary environmental information in the observed images that matches both the sub-objective and the goal of the current action. Otherwise, the current action is kept executing until encountering execution failures or the end of the episode. The successful action sequence about one sub-objective will be stored in the Performer Memory, together with necessary situational information of the agent when it was planned. For more details about the planning and execution process, please check Sup. G.2 and Sup. B.2.
@@ 3.4 Implementation Details
Percipient. The network of Percipient is depicted in Fig. 4. Images are processed by a frozen vision encoder MineCLIP [9], whose features are projected by an Alignment Net(we use two-layer MLP like LLaVA-1.5 [20]) to the same feature space as the text embeddings of the applied LLM (we use Vicuna-13B-v1.5 [5]). Then the vision and text tokens are concatenated to feed into a LoRA-based fine-tuned LLM [13]. We add LoRA [13] parameters to all projection layers of the self-attention layers in the LLM. Only the parameters of the Alignment Net and the LoRA module are optimized during training. The construction of the training data with respect to Percipient is in the Sup. B.1.
Parser, Planner, and Patroller. We utilize OpenAI’s GPT-4 [26] as LLMs in Parser, Patroller, and Planner. We also evaluate other alternatives of GPT-4 [26], such as open-source models like Vicuna-13B-v1.5 [5] and LLaMA2-70B-Chat [33] in Sup.D.3.
Performer. It is important to clarify that the actions generated by Planner are not low-level commands such as keyboard and mouse operations [1], but a set of simple actions (such as equip, move, craft). Inspired by GITM [43], we implement these actions appropriately through basic operations provided by the MineDojo [9] simulator. For more details, please check the Sup. B.2.
@@ 4.1 Experimental Setup
Environment Setting. We employ MineDojo [9] as our simulation environment to build and validate MP5. We capture player ego-view images provided by MineDojo [9] as input of MP5, and further construct a dataset for training MineLLM. As for the output of MP5, we encapsulate MineDojo’s [9] actions to create our own action space.
Task Setting. To evaluate how our MP5 can integrate perception information with planning and execution, we define two types of tasks: Context-Dependent Tasks and Process-Dependent Tasks as illustrated in Tab. 1 and Tab. 2.
1) Context-Dependent Tasks primarily study how Active Perception enables the agent to better perceive low-level context information in the environment. We first establish $6$ aspects of environmental information derived from the Minecraft game environment: [Object, Mob, Ecology, Time, Weather, Brightness]. Each aspect has multiple options. For example, pigs [Uncaptioned image], cows [Uncaptioned image], and sheep [Uncaptioned image] are all elements belonging to Mob. Based on this, we define $16$ tasks and organize their difficulty into four levels by taking into account the number of information elements that require perception, as is shown in Tab. 1. For example, Easy tasks necessitate the perception of only one element, whereas Complex tasks involve the perception of $4$ to $6$ elements. We rigorously assess MP5’s proficiency in environmental context perception across these $16$ tasks. In Context-Dependent Tasks, our environment details are predetermined (e.g., biomes [Uncaptioned image], weather [Uncaptioned image], and etc.), as certain targets are exclusive to specific environments. Without this environmental specificity, the agent might never encounter the intended target. We retain each observation of active perception throughout the task, using them as references to ascertain the agent’s successful completion of the task.
2) Process-Dependent Tasks focus on exploring the contributions of situation-aware planning, embodied action execution, and the integration with Active Perception in accomplishing long-term tasks while constantly perceiving the environment and dynamically adjusting actions. We select $25$ tasks from the technology tree and define their difficulty levels as Basic level [Uncaptioned image] to Diamond level [Uncaptioned image] based on the number of reasoning steps required to complete the tasks. All environmental factors (e.g., biomes [Uncaptioned image], weather [Uncaptioned image], and etc.) are randomized in Process-Dependent Tasks. More details can be found in Sup.D.1.
@@
Evaluation Metrics. For different tasks, the agent’s initial position and environment seed are randomized. The agent begins in survival mode, commencing with an empty inventory, and faces the challenge of hostile mob generation. It starts from scratch, with a game time limit of 10 minutes, a time period equivalent to 12,000 steps at a control frequency of 20Hz. More details can be found in Sup. C.
@@
In practice, we conduct $50$ games on Context-Dependent Tasks and $30$ games on Process-Dependent Tasks, averaging the success rates for both. The results are grouped according to the previously defined difficulty levels, and report the group means. For detailed definitions of the evaluation, please refer to Sup. D.
@@ 4.2.1 Results of Context-Dependent Tasks
In Context-Dependent Tasks, we primarily investigate how to enhance an agent’s perception of context information within the environment. We demonstrate the performance difference between Active Perception and other perception methods. We compare them with pre-trained multi-modal large language models LLaVA-1.5 [20] and GPT-4V [25], and analyze the performance of both active and fine-grained global perception on the tasks in Tab. 3. Although fine-grained global perception can obtain comprehensive perceptual information, due to the lack of objective-conditioned attention, the objective-related information obtained may be lacking or incorrect. Active perception only focuses on objective-related information and ignores other useless information, so that more accurate objective-related information can be obtained and better performance in Context-Dependent Tasks can be achieved. For the comparison, we use MineLLM, which is fine-tuned on the Minecraft instruction dataset we collect, slightly better than GPT-4V [25], which is trained on massive data, and substantially better than LLaVA-1.5 [20], which is not fine-tuned on instruction data. The complete results of Context-Dependent Tasks can be found in Sup.D.2.
@@ 4.2.2 Results of Process-Dependent Tasks
In Process-Dependent Tasks, we report the performance of the agent in completing long-horizon tasks by continuously perceiving the environment context and dynamically adjusting its actions. We also investigate the agent’s behavior in scenarios of non-situation-aware planning and non-embodied action execution. The complete results of Process-Dependent Tasks can be found in Sup.D.2.
@@ 4.2.3 Open-Ended Tasks
Processing long-horizon reasoning and understanding complex contexts are interconnected in the real world. For simplicity and comparability of the experimental setup, the first two task settings do not consider the intersection of process and context, as we cannot exhaust all combinations that these two task dimensions can form. Therefore, we refer to tasks that incorporate both Process-Dependent and Context-Dependent elements as Open-Ended Tasks. Specifically, these tasks require the agent to perceive different information of the environment at multiple stages of completing sub-objectives. As shown in Fig. 5, we present an example of an Open-Ended Task, named “Dig a block of sand [Uncaptioned image] near the water [Uncaptioned image] at night [Uncaptioned image] with a wooden shovel [Uncaptioned image]. We conduct extensive validations on this type of task, proving that MP5 can complete long-sequential tasks in challenging environments. More demonstrations and experimental results of Open-Ended Tasks can be found in Sup.F.3.